(CO 3) Sixty-five percent of products come off the line ready to ship to distributors. Your quality control department selects 12 products randomly from the line each hour. Looking at the binomial distribution, if fewer than how many are within specifications would require that the production line be shut down (unusual) and repaired?

Fewer than 4

Fewer than 5

Fewer than 10

Fewer than 6
Answer #1

Answer:

Given,

sample n = 12

p = 0.65

Mean = np

= 12*0.65

= 7.8

Standard deviation = sqrt(npq)

= sqrt(12*0.65*(1-0.65))

= sqrt(12*0.65*0.35)

= sqrt(2.73)

= 1.65

Minimum usual value = mean + 2*standard deviation

= 7.8 - 2*1.65

= 7.8 - 3.3

= 4.5

Fewer than 5 are within specifications would require that the production line be shut down (unusual) and repaired.

Option B is right answer.